From GlueXWiki
Jump to: navigation, search

October 16, 2014 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. DAQ (Beni)
  2. Electronics (Fernando)
  3. CDC update (Mike)
  4. FDC update (Vlad, Lubomir)
  5. Commissioning Wiki Page (all)
  6. Other


Participants: Curtis, Naomi (CMU), Beni, Dave, Mike, Eugene, Chris, Nick, Simon, Fernando, Vlad, Cody, Sergey, and Lubomir (JLab). .

DAQ and electronics

- Info from Beni and Cody: JTAG cable was connected (as proposed by Sergey) to the last module of crate FDC9 and chipscope installed on one of the FPGAs (the processor one). Thus, they were able to capture the problem and analyze the signals when the module didn't report data. So far, one thing is sure, the module does receive a trigger but at some point one FPGA doesn't transmit data to the next one. The plan is now to install chipscope on all FPGAs. After the problem appears even the reset doesn't help indicating is a hardware problem.

- After the meeting Sergey and Cody informed me that the problem is that the 80MHz clock sometimes doesn't come to one of the FPGAs, indicating most likely bad soldering that changes with the temperature. The plan now is to reduce the board frequency to 40MHz.

- Nick: One CAEN HV module in the CDC has been replaced. In the moment we have 3 spares. The hardware limits for the modules should be written somewhere on wiki, to be used when replacing modules (actually they are recorded in the archiver).

- Dave is implementing reading temperature and humidity in the Hall, also transmitting the FDC pressure differences as a new type epics variable.

- The gas panel in the Hall was connected to UPS.

CDC update

- Mike is working on the efficiency plots, also charge vs track length, to be used for gain calibration.

- The Bertan modules (supplying HV to two cards) will be left ON before the start of the beam alignment with a thresholds of 1 uA.

FDC update

- Vald made presentation (attached) about the strip gain calibration. You use the fact that the sum of the charge on the down strips is the same as on the upper strips. Another more important fact is that we know the shape of the charge distribution over the strips. Vlad presented results for the calibration coefficients of the third package as tested in EEL126. The gain varies by about +/-15%, smaller fraction of the variation should come from the electronics (Fernando).


- Simon and Lubomir presented their proposals for monitoring histograms. After some discussions, we agreed to have profiles of the wires and strips for each cell and package. Possibly, later, we can put the same information in 2D plots. We will also have the drift time spectra for each cell.